OpenStudy (anonymous):

Solve the following quadratic equation by completing the square, showing ALL work: 2x^2 + 12x – 4 = 0

OpenStudy (radar):

Divide thru by 2 in order to have a coefficient of 1 for the X^2. this results in: x^2 + 6x -2 = 0 Now take 1/2 of the coefficient of the x term 6 getting three and square that getting 9 now add that to both sides getting: x^2 +6x +9 - 4=9, now it is down hill the rest of the way just algebra. (X+3)^2=13 Take the square root of both sides getting: x+3=sqrt 13 \[x=-3\pm \sqrt{13}\] That is all there is to it.